OpenStudy (anonymous):

hi, please help!!! due tomorrow how do I solve this... y=2/3+4, domain(-3

OpenStudy (anonymous):

i think you typed it in wrong

OpenStudy (anonymous):

sorry, let me rewrite

OpenStudy (anonymous):

np take your time

OpenStudy (anonymous):

y=2/3x+4 domain={x: -3<x<9} not exactly sure what this is but i think its functions and relations....

OpenStudy (anonymous):

it's a equation of straight line y=mx+c; so solve it as required if you need y domain than{y: 2<y<10 }

OpenStudy (anonymous):

It is functions and relations. Sketching will show you clearly  The range of y is 2<y<10